(5a) Suppose you obtained an experimental (actual) yield of 1.81g of Aspirin and the theretically expected yield was 2.14 g. What is the % yield of Aspirin?

(5b) If your aspirin above was contaminated with 16.1% of salicylic acid, what is the correct experimental yield?

(5c) What is the correct % yield?
1
Expert's answer
2012-09-25T11:29:36-0400
For pure aspirin:
yield,% = 1.81/2.14 = 0.846 or 84.6%
For contaminated
aspirin:
yield = 1.81*(1-0.161)=1.52 g
yield,% = 1.52/2.14 = 0.710 or
71.0%